# Text Clustering with LLM Embeddings

### v4

@@ 1 Introduction
Text clustering has garnered significant interest in the field of text analysis due to its potential to uncover hidden structures within large sets of unstructured textual data. With the exponential growth of digital text content generated through platforms such as social media, online news outlets, and academic publications, the ability to organise and analyse these data has become increasingly critical. Text clustering can organise large volumes of unstructured data into meaningful categories, enabling efficient information retrieval and insightful thematic analysis across various domains, such as customer feedback, academic research, and social media content.
@@ 2.1 Text Embeddings
The landscape of text representation in natural language processing (NLP) has undergone an impressive transformation over the past few decades. From simplistic representations to highly sophisticated embeddings, advancements in this domain have significantly enhanced the ability of machines to process and understand human language with increasing accuracy.
@@ 2.2 Text Clustering Algorithms
In addition to these classic approaches, recent years have seen the rise of alternative methods for text clustering that leverage the strengths of modern embeddings and consider the unique properties of textual data. These include using deep learning models, particularly those based on autoencoders, to learn meaningful low-dimensional representations ideal for clustering [11, 12]. Ensemble clustering approaches, where multiple clustering algorithms are combined to improve the robustness and quality of the results, have also been gaining attention. These methods benefit from the diversity of the individual algorithms and can sometimes overcome the limitations of any single method [13].
@@ 3 Methods
    Preprocessing of datasets, including the removal of miscellaneous characters, emails, HTML tags, etc.
@@ 3.3 Text Embeddings
We also employed “text-embedding-ada-002” [26] embeddings, as they demonstrated the best results among OpenAI’s embeddings on larger datasets for tasks such as text search, code search, and sentence similarity. Additionally, other LLM embeddings were computed for Falcon [27] and LLaMA-2-chat [28] models, included for their respective advancements in performance and efficiency. Falcon embeddings were trained on a hybrid corpus consisting of both text documents and code, while the LLaMA-2-chat embeddings—built on the foundation of the LLaMA-2 model using an optimised auto-regressive transformer—underwent targeted fine-tuning for dialogue and question-and-answer tasks.
@@
Table 2: Tested embeddings and their configuration. For TF-IDF, parameters $\mathsf{min\_df}$ and $\mathsf{max\_df}$ are used to exclude terms that have a document frequency strictly lower than or higher than the specified thresholds, respectively, during the vocabulary creation process; if real-valued, these parameters represent a proportion of documents; if integer-valued, they represent absolute document counts. In turn, $\mathsf{max\_features}$ limit the vocabulary to only the top $\mathsf{max\_features}$ terms, ordered by term frequency across the corpus.
@@ 3.6.2 Increasing Model Dimension
Table 7: Embeddings and corresponding models used in the dimensionality experiments. Here, ‘Size’ represents the number of parameters denoted in billions (bp), and ‘Tokens’ refer to the embedding token size in trillions (trl).
@@ 4 Results and Discussion
Experiments with the MN-DS dataset, which includes a hierarchical label structure, show that clustering at a higher, more abstract label level (17 classes) yields slightly better class separation. However, the $F1$-score is higher when clustering is assessed at the more specific level (109 classes). This likely indicates better precision and recall for individual classes, while lower values for other metrics reflect the increased difficulty in maintaining overall clustering quality and cohesion with a higher number of classes. These results indicate that clustering at higher levels can produce more cohesive and interpretable clusters that align with natural categorical divisions, though at the expense of extracting less specific information for each document.
Table 8: Results of text clustering for the best-performing clustering algorithms for each combination of dataset and embedding. The best algorithm was determined by choosing the algorithm with the highest $F1$-score value. DS1 represents the CSTR dataset, DS2 is the SyskillWebert dataset, DS3 is the 20newsgroups dataset, DS4 is the MNDS dataset for level 1 labels, DS5 is the MNDS dataset for level 2 labels, and DS6 is the Reuters dataset. ‘Total’ represents the number of metrics per given embeddings/algorithm combination that outperform other combinations.
@@
Results for the model size experiment, presented in Table 10, highlight the negative influence of the number of parameters on clustering outcomes for Falcon, given that the larger model, Falcon-40b, produced lower ARI, $F1$-score, and HS for the CSTR and SyskillWebert datasets. This may be due to the fact that embeddings for the Falcon-40b model were created over a subset of the data used for the Falcon-7b embeddings (1.5 trillion tokens for Falcon-7b and 1 trillion tokens for Falcon-40b). In the case of LLaMA-2, models with sizes 7b and 13b were reportedly trained on identical datasets [41]. Results indicate that embeddings from the larger model, LLaMA-2-13b, outperform those from LLaMA-2-7b. This outcome is anticipated, as larger models with more parameters generally have a greater capacity to capture complex patterns and relationships in the data, leading to richer and more expressive embeddings.
@@ 5 Limitations
Moreover, embeddings computed for extra-large models like Falcon-180b or LLaMA-2-70b could potentially deliver substantial gains. However, practical application is restrained by the significant computational demand that accompanies increased model size. This limitation potentially skews the understanding of the absolute efficacy of the embeddings computed for these models, as performance improvements could only be inferred up to a specific size.
@@ 6 Conclusions
This research also revealed a trade-off between improved clustering results and the computational cost of embeddings from larger models. Although results indicate that an increase in model size may yield superior clustering performance, potential benefits must be weighed against the practicality of available computing resources.
@@ Acknowledgements
This work was financed by the Portuguese Agency FCT (Fundação para a Ciência e a Tecnologia), in the framework of projects UIDB/04111/2020, UIDB/00066/2020, CEECINST/00002/2021/CP2788/CT0001 and
CEECINST/00147/2018/CP1498/CT0015.

### v5

# Text Clustering with Large Language Model Embeddings

@@ Abstract
{textblock*}
190mm(-3cm,-5cm) The peer-reviewed version of this paper is published in the International Journal of Cognitive Computing in Engineering at https://doi.org/10.1016/j.ijcce.2024.11.004. This version is typeset by the authors and differs only in pagination and typographical detail.
@@ 1 Introduction
Text clustering has attracted considerable interest in text analysis due to its potential to reveal hidden structures in large volumes of unstructured textual data. With the exponential growth of digital text content generated through platforms such as social media, online news outlets, and academic publications, the ability to organise and analyse these data has become increasingly critical. Text clustering can organise large volumes of unstructured data into meaningful categories, facilitating efficient information retrieval and insightful thematic analysis across various domains, such as customer feedback, academic research, and social media content.
@@ 2.1 Text Embeddings
The field of text representation in natural language processing (NLP) has undergone an impressive transformation over the past few decades. From simple representations to highly sophisticated embeddings, advancements in this domain have significantly improved the ability of machines to process and understand human language with increasing accuracy.
@@ 2.2 Text Clustering Algorithms
In addition to these classic approaches, recent years have seen the rise of alternative methods for text clustering that leverage the strengths of modern embeddings and consider the unique properties of textual data. These include using deep learning models, particularly those based on autoencoders, to learn meaningful low-dimensional representations ideal for clustering [11, 12].
Ensemble clustering approaches, where multiple clustering algorithms are combined to improve the robustness and quality of the results, have also been gaining attention. These methods benefit from the diversity of the individual algorithms and can sometimes overcome the limitations of any single method [13].
@@ 3 Methods
    Preprocessing of datasets, including the removal of miscellaneous characters such as emails and HTML tags.
@@ 3.3 Text Embeddings
We also employed “text-embedding-ada-002” [26] embeddings, as they demonstrated the best results among OpenAI’s embeddings on larger datasets for tasks such as text search, code search, and sentence similarity. Additionally, other LLM embeddings were extracted for Falcon [27] and LLaMA-2-chat [28] models, included for their respective advancements in performance and efficiency. Falcon embeddings were trained on a hybrid corpus consisting of both text documents and code, while the LLaMA-2-chat embeddings—built on the foundation of the LLaMA-2 model using an optimised auto-regressive transformer—underwent targeted fine-tuning for dialogue and question-and-answer tasks.
@@
Table 2: Tested embeddings and their configuration. For TF-IDF, parameters $\mathsf{min\_df}$ and $\mathsf{max\_df}$ are used to exclude terms that have a document frequency strictly lower than or higher than the specified thresholds, respectively, during the vocabulary creation process; if real-valued, these parameters represent a proportion of documents; if integer-valued, they represent absolute document counts. In turn, $\mathsf{max\_features}$ limits the vocabulary to only the top $\mathsf{max\_features}$ terms, ordered by term frequency across the corpus.
@@ 3.6.2 Increasing Model Dimension
Table 7: Embeddings and corresponding models used in the dimensionality experiments. Here, ‘Size’ represents the number of parameters denoted in billions (bp), and ‘Tokens’ refers to the embedding token size in trillions (trl).
@@ 4 Results and Discussion
Experiments with the MN-DS dataset, featuring a hierarchical label structure, indicate that clustering at a higher, more abstract label level (17 classes) produces better class separation. However, the $F1$-score is higher when clustering is assessed at the more specific level (109 classes). This likely indicates better precision and recall for individual classes, while lower values for other metrics reflect the increased difficulty in maintaining overall clustering quality and cohesion with a higher number of classes. These results indicate that clustering at higher levels can produce more cohesive and interpretable clusters that align with natural categorical divisions, though at the expense of extracting less specific information for each document.
Table 8: Results of text clustering for the best-performing clustering algorithms for each combination of dataset and embedding. The best algorithm was determined by choosing the algorithm with the highest $F1$-score value. DS1 represents the CSTR dataset, DS2 is the SyskillWebert dataset, DS3 is the 20newsgroups dataset, DS4 is the MN-DS dataset for level 1 labels, DS5 is the MN-DS dataset for level 2 labels, and DS6 is the Reuters dataset. ‘Total’ represents the number of metrics per given embeddings/algorithm combination that outperform other combinations.
@@
Results for the model size experiment, presented in Table 10, highlight the negative influence of the number of parameters on clustering outcomes for Falcon, given that the larger model, Falcon-40b, produced lower ARI, $F1$-score, and HS for the CSTR and SyskillWebert datasets. This may be because embeddings for the Falcon-40b model were created over a subset of the data used for the Falcon-7b embeddings (1.5 trillion tokens for Falcon-7b and 1 trillion tokens for Falcon-40b). In the case of LLaMA-2, models with sizes 7b and 13b were reportedly trained on identical datasets [41]. Results indicate that embeddings from the larger model, LLaMA-2-13b, outperform those from LLaMA-2-7b. This outcome is anticipated, as larger models with more parameters generally have a greater capacity to capture complex patterns and relationships in the data, leading to richer and more expressive embeddings.
@@ 5 Limitations
Moreover, embeddings computed for extra-large models like Falcon-180b or LLaMA-2-70b could potentially deliver substantial gains. However, practical application is restrained by the significant computational demands associated with larger model sizes. This limitation potentially skews the understanding of the absolute efficacy of the embeddings computed for these models, as performance improvements could only be inferred up to a specific size.
@@ 6 Conclusions
This research highlights the trade-off between improved clustering performance and the computational costs of using embeddings from larger models. Although results indicate that an increase in model size may yield superior clustering performance, potential benefits must be weighed against the practicality of available computing resources.
@@ Acknowledgements
This work was financed by the Fundação para a Ciência e a Tecnologia, in the framework of projects UIDB/04111/2020, UIDB/00066/2020, CEECINST/00002/2021/CP2788/CT0001, CEECINST/00147/2018/CP1498/CT0015 as well as Instituto Lusófono de Investigação e Desenvolvimento (ILIND) under project COFAC/ILIND/COPELABS/1/2022.